OpenStudy (anonymous):

Use chain rule to solve y=sin^2x

zepdrix (zepdrix):

\[\Large\rm \sin^2x=(\sin x)^2\]It might be easier to work with in this form.

zepdrix (zepdrix):

So... power rule to start, yes?

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\left[(\sin x)^2\right]'}=2(\sin x)\color{royalblue}{\left[\sin x\right]'}\]Power rule then we chain.

OpenStudy (anonymous):

which would be 2(sinx)*(cos)?

zepdrix (zepdrix):

Mmmm yah looks good!

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\left[(\sin x)^2\right]'}=2(\sin x)\color{orangered}{\left[\cos x\right]}\]

OpenStudy (anonymous):

Thank You! :D